# Almost real closed fields with real analytic structure

### v1

@@ 4.2. Eliminating to the residue field and value group
Let ${\mathcal{B}}$ be a real Weierstrass system over $\Omega$ which is rich, and let $K$ be an almost real closed field with ${\mathcal{B}}$-analytic structure. Then $K$ equipped with the canonical valuation has ${\mathcal{L}}_{{\rm ac},{\mathcal{B}}}$-structure such that ${\rm ac}$ and ${\mathcal{B}}$ are compatible.
@@
For an arbitrary convex valuation, one can still obtain relative quantifier elimination by adding symbols for all elements of ${\mathcal{B}}$ to the language on the residue field. However, it is then not so clear what the language on the residue field actually is. We expect that in general, the residue field will be another almost real closed field equipped with ${\mathcal{B}}^{\prime\prime}$-analytic structure, for some related Weierstrass system ${\mathcal{B}}^{\prime\prime}$ over a quotient $\Omega^{\prime\prime}$ of $\Omega$. In fact, this is what happens above for the canonical valuation. Indeed, the new Weierstrass system ${\mathcal{B}}^{\prime}$ constructed above has value group $\Omega^{\prime}=0$, which is related to the fact that the residue field ${\mathbb{R}}$ has value group 00.
@@ 5.1. Hensel minimality
    (Jacobian property [3, Lem. 2.8.5]) If $f:K\to K$ is a definable function, then there exists a finite set $C\subset K$ such that for each ball $B$ disjoint from $C$ there exists a $\lambda\in\Gamma$ such that for all $x,y\in B$
@@
    (Differentiation [3, Cor. 3.1.6]) Continuing with the notation from the Jacobian property, there exists a finite set $C\subset K$ such that on each ball $B$ disjoint from $C$, $f$ is $C^{1}$, $v(f^{\prime}(x))$ is constant for $x\in B$, and for all $x,y\in B$
@@
    (Cell decomposition [3, Thm. 5.2.4]) There exists a good notion of cells such that every definable set $X\subset K^{n}$ may be written as a finite union of definable cells. Using cells, one also obtains a good dimension theory of definable sets [3, Prop. 5.3.4].
@@
    (Resplendency [3, Thm. 4.1.19]) If ${\mathcal{L}}^{\prime}\supset{\mathcal{L}}_{{\mathrm{val}},{\mathcal{B}}}$ is any expansion of the language by predicates on Cartesian powers of $\mathrm{RV}$, then $\operatorname{Th}_{{\mathcal{L}}^{\prime}}(K)$ is also $\omega$-h-minimal. For example, we may add an angular component map ${\rm ac}:K\to k$ by adding a splitting $\Gamma\to\mathrm{RV}$, and the resulting theory will still be $\omega$-h-minimal.
@@
We use the notion of rational points and counting dimension from [2]. Consider $K={\mathbb{R}}((t))$ for the canonical valuation and equipped with analytic structure from the full Weierstrass system ${\mathcal{A}}(({\mathbb{Z}}))$. Let $X$ be the graph of the exponential function $\exp:[-1,1]_{K}\to K$, which is a definable transcendental curve. For $s\geq 1$ an integer, let $X_{s}$ denote the rational points of height at most $s$, i.e. those points $(x,y)\in X$ for which $x,y\in{\mathbb{R}}[t]$ are polynomials of degree at most $s-1$. Then $X_{1}$ is infinite, as it contains all points of the form $(x,\exp(x))$ for $x\in{\mathbb{R}}$. However, for every $s\geq 1$ we have that $X_{s}=X_{1}$, and so the counting dimension of $X$ is bounded by $(1,1,1)$.
In general, if $X\subset K^{n}$ is a definable transcendental curve, then [2, Thm. 2.2.1] shows that for every $\varepsilon>0$, the counting dimension of $X$ is bounded by $(N_{\varepsilon},1,\lceil\varepsilon s\rceil)$, where $N_{\varepsilon}$ is a constant depending on $\varepsilon$.
@@ 5.2. The induced structure on the residue field and value group
Let ${\mathcal{B}}$ be a real Weierstrass system which is strong and rich. Let $K$ be an almost real closed field equipped with ${\mathcal{B}}$-analytic structure, considered as a valued field for the canonical valuation, and equip it with ${\mathcal{L}}_{{\rm ac},{\mathcal{B}}}$-structure. Assume that ${\rm ac}$ and ${\mathcal{B}}$ are compatible, then
@@
The first part follows from resplendency [3, Thm. 4.1.19], since adding an ${\rm ac}$ map to ${\mathcal{L}}_{{\mathrm{val}},{\mathcal{B}}}$ is an $\mathrm{RV}$-expansion. The other parts follow directly from the relative quantifier elimination result from Proposition 4.2.3. ∎
@@ 5.3. An Ax–Kochen–Ersov theorem
Let ${\mathcal{B}}$ be a real Weierstrass system which is strong and rich. Let $K$ and $K^{\prime}$ be almost real closed fields equipped with ${\mathcal{B}}$-analytic structure, considered as valued fields for the canonical valuation, and equip them with ${\mathcal{L}}_{{\rm ac},{\mathcal{B}}}$-structure. Let $G$ and $G^{\prime}$ be the value groups of $K$ and $K^{\prime}$ and assume that ${\rm ac}$ and ${\mathcal{B}}$ are compatible. Then

### v2

@@ 4.2. Eliminating to the residue field and value group
Let ${\mathcal{B}}$ be a real Weierstrass system over $\Omega$ which is rich, and let $K$ be an almost real closed field with ${\mathcal{B}}$-analytic structure. Then $K$ equipped with the natural valuation has ${\mathcal{L}}_{{\rm ac},{\mathcal{B}}}$-structure such that ${\rm ac}$ and ${\mathcal{B}}$ are compatible.
@@
For an arbitrary convex valuation, one can still obtain relative quantifier elimination by adding symbols for all elements of ${\mathcal{B}}$ to the language on the residue field. However, it is then not so clear what the language on the residue field actually is. We expect that in general, the residue field will be another almost real closed field equipped with ${\mathcal{B}}^{\prime\prime}$-analytic structure, for some related Weierstrass system ${\mathcal{B}}^{\prime\prime}$ over a quotient $\Omega^{\prime\prime}$ of $\Omega$. In fact, this is what happens above for the natural valuation. Indeed, the new Weierstrass system ${\mathcal{B}}^{\prime}$ constructed above has value group $\Omega^{\prime}=0$, which is related to the fact that the residue field ${\mathbb{R}}$ has value group 00.
@@ 5.1. Hensel minimality
    (Jacobian property [6, Lem. 2.8.5]) If $f:K\to K$ is a definable function, then there exists a finite set $C\subset K$ such that for each ball $B$ disjoint from $C$ there exists a $\lambda\in\Gamma$ such that for all $x,y\in B$
@@
    (Differentiation [6, Cor. 3.1.6]) Continuing with the notation from the Jacobian property, there exists a finite set $C\subset K$ such that on each ball $B$ disjoint from $C$, $f$ is $C^{1}$, $v(f^{\prime}(x))$ is constant for $x\in B$, and for all $x,y\in B$
@@
    (Cell decomposition [6, Thm. 5.2.4]) There exists a good notion of cells such that every definable set $X\subset K^{n}$ may be written as a finite union of definable cells. Using cells, one also obtains a good dimension theory of definable sets [6, Prop. 5.3.4].
@@
    (Resplendency [6, Thm. 4.1.19]) If ${\mathcal{L}}^{\prime}\supset{\mathcal{L}}_{{\mathrm{val}},{\mathcal{B}}}$ is any expansion of the language by predicates on Cartesian powers of $\mathrm{RV}$, then $\operatorname{Th}_{{\mathcal{L}}^{\prime}}(K)$ is also $\omega$-h-minimal. For example, we may add an angular component map ${\rm ac}:K\to k$ by adding a splitting $\Gamma\to\mathrm{RV}$, and the resulting theory will still be $\omega$-h-minimal.
@@
We use the notion of rational points and counting dimension from [5]. Consider $K={\mathbb{R}}((t))$ for the natural valuation and equipped with analytic structure from the full Weierstrass system ${\mathcal{A}}(({\mathbb{Z}}))$. Let $X$ be the graph of the exponential function $\exp:[-1,1]_{K}\to K$, which is a definable transcendental curve. For $s\geq 1$ an integer, let $X_{s}$ denote the rational points of height at most $s$, i.e. those points $(x,y)\in X$ for which $x,y\in{\mathbb{R}}[t]$ are polynomials of degree at most $s-1$. Then $X_{1}$ is infinite, as it contains all points of the form $(x,\exp(x))$ for $x\in{\mathbb{R}}$. However, for every $s\geq 1$ we have that $X_{s}=X_{1}$, and so the counting dimension of $X$ is bounded by $(1,1,1)$.
In general, if $X\subset K^{n}$ is a definable transcendental curve, then [5, Thm. 2.2.1] shows that for every $\varepsilon>0$, the counting dimension of $X$ is bounded by $(N_{\varepsilon},1,\lceil\varepsilon s\rceil)$, where $N_{\varepsilon}$ is a constant depending on $\varepsilon$.
Additionally, $1$-h-minimal fields automatically satisfy many other nice topological properties, by [29, 31, 30]. Note that the statements [31] require that the residue field and value group are orthogonal, and that the structure on the latter is the pure ordered abelian group structure. By Theorem 1.0.2, our ordered fields with real analytic structure do indeed satisfy this additional condition.
@@ 5.2. The induced structure on the residue field and value group
Let ${\mathcal{B}}$ be a real Weierstrass system which is strong and rich. Let $K$ be an almost real closed field equipped with ${\mathcal{B}}$-analytic structure, considered as a valued field for the natural valuation, and equip it with ${\mathcal{L}}_{{\rm ac},{\mathcal{B}}}$-structure. Assume that ${\rm ac}$ and ${\mathcal{B}}$ are compatible, then
@@
The first part follows from resplendency [6, Thm. 4.1.19], since adding an ${\rm ac}$ map to ${\mathcal{L}}_{{\mathrm{val}},{\mathcal{B}}}$ is an $\mathrm{RV}$-expansion. The other parts follow directly from the relative quantifier elimination result from Proposition 4.2.3. ∎
@@
5.2.2 Remark.
The above result is complementary to the main Theorem by Cubides-Kovacsics and Haskell in [12]. They investigate real closed fields with separated and overconvergent analytic structure (instead of real analytic structure), and prove quantifier elimination in a one-sorted language. From this they deduce that such structures are weakly o-minimal.
In our setting, an appropriate version of the latter also holds: if $K$ is a real closed field with ${\mathcal{B}}$-analytic structure for some real Weierstrass system ${\mathcal{B}}$, then $T_{{\mathcal{L}}_{{\mathcal{B}}}}(K)$ is weakly o-minimal. Indeed, as $T_{{\mathcal{L}}_{{\mathcal{B}}}}(K)$ is the expansion of an o-minimal structure by a convex predicate, it is weakly o-minimal by [1, § 4] (see also [2]).
@@ 5.3. An Ax–Kochen–Ersov theorem
Let ${\mathcal{B}}$ be a real Weierstrass system which is strong and rich. Let $K$ and $K^{\prime}$ be almost real closed fields equipped with ${\mathcal{B}}$-analytic structure, considered as valued fields for the natural valuation, and equip them with ${\mathcal{L}}_{{\rm ac},{\mathcal{B}}}$-structure. Let $G$ and $G^{\prime}$ be the value groups of $K$ and $K^{\prime}$ and assume that ${\rm ac}$ and ${\mathcal{B}}$ are compatible. Then